OpenStudy (superflytyguy):

Simplify the trigonometric expression. Show your work. Really need help on this one

OpenStudy (superflytyguy):

\[ \frac{ 1 }{ 1 +\sin \theta } + \frac{ 1 }{ 1 - \sin \theta }\]

OpenStudy (highschoolmom2010):

\[\frac{ 1- \sin \theta }{ (1+ \sin \theta)(1-\sin \theta) }+\frac{ 1+\sin \theta }{ (1+\sin \theta)(1-\sin \theta) }\] \[\frac{ 1-\sin \theta +1+\sin \theta }{ (1+\sin \theta)(1-\sin \theta) }\] \[\frac{ 2 }{ 1- \sin^2 \theta }\] \[\frac{ 2 }{ \cos^2 \theta }\] \[2 \sec ^2 \theta\]

OpenStudy (superflytyguy):

Thank you so much!!!

OpenStudy (highschoolmom2010):

no problem :)